OpenStudy (anonymous):

Help! I'm not sure how to solve this. Ms. Lunette has 18 students in her class. She wants to send 3 of these students to pick up books for the class. How many combinations of 3 students can she choose? A. 6 B. 54 C. 816 D. 4896

OpenStudy (anonymous):

She has 18 and wants to choose 3

OpenStudy (mathstudent55):

This is a problem in which order does not matter. Let's say the 18 students are named 1 through 18 If she picks 1, 2, 3, that's the same as picking 3, 2, 1. Once again, order does not matter.

OpenStudy (anonymous):

It is hard to visualize, need the formula

OpenStudy (anonymous):

ok then. like this? 123. 456. 789. 101112. 131415. 161718.

OpenStudy (mathstudent55):

This is a combination problem (not permutation). \(_nC_r = \dfrac{n!}{r!(n - r)!} \) \(_{18}C_3 = \dfrac{18!}{3!(18 - 3)!}\)

OpenStudy (anonymous):

I am so bad at this. I think it's c

OpenStudy (anonymous):

i mean A

OpenStudy (mathstudent55):

How about just doing it together and finding out?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

 Ok so far?